# Multipoint Schwarz–Pick Lemma for the quaternionic case

### v3

@@ Introduction
Over the skew-field of quaternions $\mathbb{H}$, which contains $\mathbb{C}$, the notion of slice regularity takes the place of holomorphicity. Slice regular functions have been introduced by Gentili and Struppa [17], developing a regularity concept that was originally formulated by Cullen [14]. It is based on fact that it can be given a ‘book structure’ to the skew-field of quaternions $\mathbb{H}=\bigcup_{I\in\mathbb{S}}\mathbb{C}_{I}$, where $\mathbb{S}$ is the sphere of imaginary units $\{q\in\mathbb{H}:q^{2}=-1\}$ and $\mathbb{C}_{I}$ are the complex subspaces $\{x+yI\in\mathbb{H}:x,y\in\mathbb{R}\}$, which are known as complex slices and which can be identified with complex planes having the real line $\mathbb{R}$ in common. A function $f:\Omega\longrightarrow\mathbb{H}$ defined on a domain in $\mathbb{H}$ is said to be slice regular if and only if its restriction to any slice $\mathbb{C}_{I}$ is holomorphic, i.e., the differential operator ¯$\overline{\partial}_{I}=1/2(\partial_{x}+I\partial_{y})$ vanishes on $f\mid_{\Omega\cap\mathbb{C}_{I}}$. If so, one can define a derivative $\partial_{C}f$ (the regular, or Cullen, derivative) in this way: $\partial_{C}f(q)=\partial_{I}(f\mid_{\Omega\cap\mathbb{C}_{I}})(q)$ for any $q\in\Omega\cap\mathbb{C}_{I}$, where $\partial_{I}$ is the operator $1/2(\partial_{x}-I\partial_{y})$. It is important to observe that the common pointwise product of functions does not preserve slice regularity; however, one could define a different product on spaces of slice regular functions, the $\ast$-product $f\ast g$, which allows to define on them the structure of a (noncommutative) algebra. Using this operation together with the $\ast$-inversion $f\mapsto f^{-\ast}$, one can define slice regular Möbius transformations on the quaternionic unit ball $\mathbb{B}=\{q:\mathbb{H}:|q|<1\}$, that is to say the maps of the form
@@
All the facts we need from the theory of slice regularity are reviewed in Section 1, with a focus on slice regular Möbius transformations (following [11, 22]) and regular Blaschke products on $\mathbb{B}$. For further information we refer to the recent monograph by Gentili, Stoppato and Struppa [16].
@@
In Section 3 we show some consequences of this results. In particular, following [7] and restricting to the case of a fixed point at the origin, by the connection between the regular derivative $\partial_{C}f(p)$ and the hyperbolic derivative $f^{h}(p):=f^{\star}_{p}(p)$ for slice regular self-maps on $\mathbb{B}$, we get the quaternionic version of some inequalities well established in classical complex analysis, such as Dieudonné’s or Goluzin’s estimates (assuming $f(0)=0$).
@@
In Section 5 we consider the Nevanlinna–Pick interpolation problem with $n$ points in the quaternionic unit ball $\mathbb{B}$, looking for a slice regular interpolating function $f:\mathbb{B}\longrightarrow\mathbb{B}$ for the prescribed nodes and values. Here we give a generalization to the quaternionic setting of the conditions and the algorithm presented in [6] for the complex case, in terms of hyperbolic difference quotients. Since the quaternionic setting is wilder than the complex one and not commutative, we have to restrict the problem to the case of real nodes. Given $n$ distinct real points $r_{1},\ldots,r_{n}\in(-1,1)$ and $n$ values $s_{1},\ldots,s_{n}$ in the quaternionic unit ball $\mathbb{B}$, one defines by iteration quantities $Q_{\kappa}^{\ell}$ for $\kappa=1,\ldots,n-1$ and $\ell=\kappa+1,\ldots,n$ which are generically expressed as quotients of Möbius maps, starting from the given nodes and values. Then we have a criterion in terms of the last term $Q_{n-1}^{n}$:
@@
When the nodes are not all real, the algorithm does not work anymore: we give some evidences in the last part of Section 5.
@@ 1.1. Slice regular functions
Let $\Omega\subseteq\mathbb{H}$ be a domain. For any $I\in\mathbb{S}$, let $\Omega_{I}:=\Omega\cap\mathbb{C}_{I}=\{x+yI\in\Omega:x,y\in\mathbb{R}\}$. A function $f:\Omega\longrightarrow\mathbb{H}$ is said to be (left) slice regular on $\Omega$ ($f\in\mathcal{SR}(\Omega)$) if it has continuous partial derivatives and any restriction $f_{I}=f\mid_{\Omega_{I}}$ for $I\in\mathbb{S}$ is holomorphic, in the sense that ¯$\overline{\partial}_{I}f_{I}(x+yI)\equiv 0$ on $\Omega_{I}$, where ¯$\overline{\partial}_{I}$ is the differential operator
¯$\overline{\partial}_{I}:=\phi_{I}\mapsto\frac{1}{2}\left(\partial_{x}\phi_{I}+I\partial_{y}\phi_{I}\right),\qquad\phi_{I}:\Omega_{I}\longrightarrow\mathbb{H}.$
@@
On the other side, the function $q\mapsto i\cdot q$ is not slice regular, since if we consider the slice $\mathbb{C}_{j}$, then the relative restriction is $x+yj\mapsto xi+yk$ and ¯$\overline{\partial}_{j}(xi+yk)=\frac{1}{2}(i+jk)=\frac{2i}{2}=i$.
@@
The sum $f+g$ of two slice regular functions $f,g\in\mathcal{SR}(\Omega)$ is still slice regular. The last example shows that the standard pointwise product $fg$ of two slice regular functions $f,g$ is not always slice regular. Moreover, the composition of two slice regular functions $g\circ f$, if defined, is not in general slice regular: for instance $q\mapsto(q\cdot i)^{2}=q\cdot i\cdot q\cdot i$ is such that its restriction to $\mathbb{C}_{j}$ is $(x+yj)\mapsto(xi-yk)^{2}=-x^{2}-y^{2}$ and ¯$\overline{\partial}_{j}(-x^{2}-y^{2})=-(x+yj)\neq 0$ for all $x+yj\neq 0$.
@@
One may also define a right regular quotient $g\ast f^{-1}$, which is still slice regular in $B(0,R)\setminus\mathcal{Z}(f^{s})$; in general it is different from $f^{-\ast}\ast g$, but they coincide when $f$ and $g$ast-commute.
@@ 1.2. Regular linear fractional and Möbius transformations
It is clear that for all $\lambda\in\mathbb{R}\setminus\{0\}$, it is $\mathcal{F}_{\lambda A}=\mathcal{F}_{A}$, hence we can consider $A$ belonging to a class in the quotient group $\operatorname{PSL}(2,\mathbb{H})=\operatorname{GL}(2,\mathbb{H})/\{\lambda\mathbf{I}:\lambda\in\mathbb{R}\setminus\{0\}\}$. In particular, we may choose $\lambda=(\det_{\mathbb{H}}A)^{-1/2}$ and so we can assume $A\in\operatorname{SL}(2,\mathbb{H})=\{A\in\operatorname{GL}(2,\mathbb{H}):\det_{\mathbb{H}}A=1\}$.
@@
Proposition 1.17 ([22, Cor. 7.2]).
Let $\mathcal{F}_{A}$ be a regular linear fractional transformation for some $A\in\operatorname{GL}(2,\mathbb{H})$. The following are equivalent:
@@
[…]
@@ 1.3. Regular Blaschke products
Let us remark that $|\gamma|=1$. Using this trick and then proceeding as in the $\kappa=2$ case, one gets that the equation $B(q)=q_{0}$ is equivalent to $P_{\kappa}(q)=0$, where $P_{\kappa}$ is a slice regular polynomial of degree $\kappa$, hence the set of its roots has total multiplicity $\kappa=n+2m$, where $m$ is number of spherical zeros and $n$ number of non-spherical zeros (counted with their respective multiplicity).
@@ 2. The hyperbolic difference quotient and the three-points Schwarz–Pick Lemma
At a first look, since the term $(q-p)^{-\ast}$ is well defined only for $q\in\mathbb{B}$ not in the sphere $\mathbb{S}_{p}=\{q:q^{2}-q\cdot 2\operatorname{Re}p+|p|^{2}=0\}$, the domain of $\mathcal{R}_{f,p}$ is only $\mathbb{B}\setminus\mathbb{S}_{p}$. But the singularity at $\mathbb{S}_{p}$ is removable: indeed, by applying [16, Prop. 3.18] to $f-f(p)$, there exists a unique slice regular function $g:\mathbb{B}\longrightarrow\mathbb{B}$ for which $f(q)=f(p)+(q-p)\ast g(q)$, and it is clear that $g$ coincides with $\mathcal{R}_{f,p}$ in $\mathbb{H}\setminus\mathbb{S}_{p}$. The value in $p$ of this function coincides with the regular derivative $\partial_{C}f(p)$, whereas its value in $\overline{p}$ coincides by the very definition with the spherical derivative in $p$: $\partial_{S}f(p):=(2\operatorname{Im}p)^{-1}(f(p)-f(\overline{p}))$.
@@
(7) $f^{h}(p)=f_{p}^{\star}(p)=(1-|p|^{2})\partial_{C}f(p)(1-\overline{f(p)}f(p_{0}))^{-1},$
for some $p_{0}\in\mathbb{S}_{p}$ depending on $p$ and on the function $f$ itself. Note that $f_{p}^{\star}$ can also be written as
$\boxed{f_{p}^{\star}=\mathcal{M}_{p}^{-\ast}\ast(\mathcal{M}_{f(p)}\bullet f)}$
@@ 2.2. Three-points quaternionic Schwarz–Pick Lemma
3. Some consequences. Estimates in the case of $\boldsymbol{f(0)=0}$
@@
Since $\mathcal{M}_{0}$ is the identity function $q\mapsto q$ and $\mathcal{M}_{0}\bullet f=f$, it is clear that $f^{\star}_{0}(q_{0})=(\mathcal{M}_{0}^{-\ast}\ast(\mathcal{M}_{0}\bullet f))(q_{0})=q_{0}^{-1}f(q_{0})$. On the other side, since 00 is invariant by the rotations appearing by evaluating the $\ast$-products punctually, then
@@
Using the fact that $\mathcal{M}_{r}=M_{r}$ is a slice preserving function and that $\mathcal{M}_{f(r)}\bullet f$ coincides with $M_{f(r)}\circ f$ on the real interval $(-1,1)$, one gets
@@
and the factor $-(1-q\overline{p})^{-1}(1-p\overline{q})$ has clearly modulus one. ∎
@@
By the triangle inequality and (12) with $p=0,q=s$ and $f$ replaced by $f^{\star}_{r}$, it is

### v4

@@ Introduction
Over the skew-field of quaternions $\mathbb{H}$, which contains $\mathbb{C}$, the notion of slice regularity takes the place of holomorphicity. Slice regular functions have been introduced by Gentili and Struppa [17], developing a regularity concept that was originally formulated by Cullen [14]. It is based on the fact that it can be given a ‘book structure’ to the skew-field of quaternions $\mathbb{H}=\bigcup_{I\in\mathbb{S}}\mathbb{C}_{I}$, where $\mathbb{S}$ is the sphere of imaginary units $\{q\in\mathbb{H}:q^{2}=-1\}$ and $\mathbb{C}_{I}$ are the complex subspaces $\{x+yI\in\mathbb{H}:x,y\in\mathbb{R}\}$, which are known as complex slices and which can be identified with complex planes having the real line $\mathbb{R}$ in common. A function $f:\Omega\longrightarrow\mathbb{H}$ defined on a domain in $\mathbb{H}$ is said to be slice regular if and only if its restriction to any slice $\mathbb{C}_{I}$ is holomorphic, i.e., the differential operator $\overline{\partial}_{I}=1/2(\partial_{x}+I\partial_{y})$ vanishes on $f\mid_{\Omega\cap\mathbb{C}_{I}}$. If so, one can define a derivative $\partial_{C}f$ (the regular, or Cullen, derivative) in this way: $\partial_{C}f(q)=\partial_{I}(f\mid_{\Omega\cap\mathbb{C}_{I}})(q)$ for any $q\in\Omega\cap\mathbb{C}_{I}$, where $\partial_{I}$ is the operator $1/2(\partial_{x}-I\partial_{y})$. It is important to observe that the common pointwise product of functions does not preserve slice regularity; however, one could define a different product on spaces of slice regular functions, the $\ast$-product $f\ast g$, which allows to define on them the structure of a (noncommutative) algebra. Using this operation together with the $\ast$-inversion $f\mapsto f^{-\ast}$, one can define slice regular Möbius transformations on the quaternionic unit ball $\mathbb{B}=\{q:\mathbb{H}:|q|<1\}$, that is to say the maps of the form
@@
All the facts we need from the theory of slice regularity are reviewed in Section 1, with a focus on slice regular Möbius transformations (following [11, 23]) and regular Blaschke products on $\mathbb{B}$. The main difference between the classical complex setting and the quaternionic frame is essentially due to the contrasts between the theory of slice regularity with respect to the classical holomorphicity, which are stressed in Section 1. We recall, for instance, the fact that we cannot use compositions of functions as in the complex case, but we need to consider an action of a matrix group on slice regular functions. These discrepancies, due essentially to the non-commutativity of the quaternions, have some consequences in trying to apply arguments from complex analysis to this novel framework: we will need to choose some points along the real line (which is incidentally the center of the skew-field of quaternions). For further information about the theory of quaternionic slice regularity we refer to the recent monograph by Gentili, Stoppato and Struppa [16].
@@
In Section 3 we show some consequences of these results. In particular, following [7] and restricting to the case of a fixed point at the origin, by the connection between the slice regular derivative $\partial_{C}f(p)$ and the hyperbolic derivative $f^{h}(p):=f^{\star}_{p}(p)$ for slice regular self-maps on $\mathbb{B}$, we get the quaternionic version of some inequalities well established in classical complex analysis, such as Dieudonné’s or Goluzin’s estimates (assuming $f(0)=0$).
@@
In Section 5 we consider the Nevanlinna–Pick interpolation problem with $n$ points in the quaternionic unit ball $\mathbb{B}$, looking for a slice regular interpolating function $f:\mathbb{B}\longrightarrow\mathbb{B}$ for the prescribed nodes and values. Here we give a generalization to the quaternionic setting of the conditions and the algorithm presented in [6] for the complex case, in terms of hyperbolic difference quotients. As recalled before, since the quaternionic setting is wilder than the complex one and not commutative, we have to restrict the problem to the case of real nodes. Given $n$ distinct real points $r_{1},\ldots,r_{n}\in(-1,1)$ and $n$ values $s_{1},\ldots,s_{n}$ in the quaternionic unit ball $\mathbb{B}$, one defines by iteration quantities $Q_{\kappa}^{\ell}$ for $\kappa=1,\ldots,n-1$ and $\ell=\kappa+1,\ldots,n$ which are generically expressed as quotients of Möbius maps, starting from the given nodes and values. Then we have a criterion in terms of the last term $Q_{n-1}^{n}$:
@@
When the nodes are not all real, the algorithm does not work anymore: we give some evidence in the last part of Section 5.
@@ 1.1. Slice regular functions
Let $\Omega\subseteq\mathbb{H}$ be a domain. For any $I\in\mathbb{S}$, let $\Omega_{I}:=\Omega\cap\mathbb{C}_{I}=\{x+yI\in\Omega:x,y\in\mathbb{R}\}$. A function $f:\Omega\longrightarrow\mathbb{H}$ is said to be (left) slice regular on $\Omega$ ($f\in\mathcal{SR}(\Omega)$) if it has continuous partial derivatives and any restriction $f_{I}=f\mid_{\Omega_{I}}$ for $I\in\mathbb{S}$ is holomorphic, in the sense that $\overline{\partial}_{I}f_{I}(x+yI)\equiv 0$ on $\Omega_{I}$, where $\overline{\partial}_{I}$ is the differential operator
$\overline{\partial}_{I}:=\phi_{I}\mapsto\frac{1}{2}\left(\partial_{x}\phi_{I}+I\partial_{y}\phi_{I}\right),\qquad\phi_{I}:\Omega_{I}\longrightarrow\mathbb{H}.$
@@
On the other side, the function $q\mapsto i\cdot q$ is not slice regular, since if we consider the slice $\mathbb{C}_{j}$, then the relative restriction is $x+yj\mapsto xi+yk$ and $\overline{\partial}_{j}(xi+yk)=\frac{1}{2}(i+jk)=\frac{2i}{2}=i$.
@@
The sum $f+g$ of two slice regular functions $f,g\in\mathcal{SR}(\Omega)$ is still slice regular. The last example shows that the standard pointwise product $fg$ of two slice regular functions $f,g$ is not always slice regular. Moreover, the composition of two slice regular functions $g\circ f$, if defined, is not in general slice regular: for instance $q\mapsto(q\cdot i)^{2}=q\cdot i\cdot q\cdot i$ is such that its restriction to $\mathbb{C}_{j}$ is $(x+yj)\mapsto(xi-yk)^{2}=-x^{2}-y^{2}$ and $\overline{\partial}_{j}(-x^{2}-y^{2})=-(x+yj)\neq 0$ for all $x+yj\neq 0$.
@@
One may also define a right regular quotient $g\ast f^{-\ast}$, which is still slice regular in $B(0,R)\setminus\mathcal{Z}(f^{s})$; in general it is different from $f^{-\ast}\ast g$, but they coincide when $f$ and $g$ $\ast$-commute.
@@ 1.2. Regular linear fractional and Möbius transformations
It is clear that for all $\lambda\in\mathbb{R}\setminus\{0\}$, it is $\mathcal{F}_{\lambda A}=\mathcal{F}_{A}$, hence we can consider $A$ belonging to a class in the quotient group $\operatorname{PSL}(2,\mathbb{H})=\operatorname{GL}(2,\mathbb{H})/\{\lambda\mathbf{I}:\lambda\in\mathbb{R}\setminus\{0\}\}$. In particular, we may choose $\lambda=(\det_{\mathbb{H}}A)^{-1/2}$ and so we can assume $A\in\operatorname{SL}(2,\mathbb{H})=\{A\in\operatorname{GL}(2,\mathbb{H}):\det_{\mathbb{H}}A=1\}$. Be aware that the reduction to $\operatorname{PSL}(2,\mathbb{H})$ is not enough to have a bijective correspondence between matrices and transformations: for instance, $\mathcal{F}_{c\mathbf{I}}(q)=\mathcal{F}_{\mathbf{I}}=\operatorname{id}$ for any $c\in\mathbb{H}\setminus\{0\}$.
@@
Proposition 1.17 ([23, Cor. 7.2]).
Let $\mathcal{F}_{A}$ be a regular linear fractional transformation associated to some $A\in\operatorname{GL}(2,\mathbb{H})$. The following are equivalent:
@@
belongs to $\mathcal{SR}(\mathbb{B},\mathbb{B})$ and the map $\operatorname{Sp}(1,1)\times\mathcal{SR(\mathbb{B},\mathbb{B})}\longrightarrow\mathcal{SR(\mathbb{B},\mathbb{B})},(A,f)\mapsto f.A$ is a right action, i.e., $(f.A).B=f.(AB)$.
@@
We will use the following notation from now on:
Definition 1.20.
If $A=(1-|p|^{2})^{-1/2}\begin{pmatrix}1&-\overline{p}\\ -p&1\end{pmatrix}\in\operatorname{Sp}(1,1)$ and $f$ is any function in $\mathcal{SR}(\mathbb{B},\mathbb{B})$, then we denote by $\mathcal{M}_{p}\bullet f$ the function
@@
Evaluating at a point $q\in\mathbb{B}$ such function gives
@@
This notation is motivated by the fact that $\mathcal{F}_{A}$ coincides with $\mathcal{M}_{p}$ for $A$ given as above, and that represents, in some sense, a regular counterpart of the composition of a Möbius map with a self-map $f$ of the unit ball. However, one should be careful: in general, we cannot replace $A$ with any matrix $B$ such that $\mathcal{F}_{B}=\mathcal{M}_{p}$: we observed in Remark 1.15 that the identity function, which coincides with $\mathcal{M}_{0}$, can be represented by any multiple $c\mathbf{I}$ of the identity matrix by a nonzero quaternion $c$, however it we choose $c=i$ and $f(q)=qj$, then it is easy to show that $f.\mathbf{I}(q)=f(q)=qj$ while $f.(i\mathbf{I})(q)=i\ast qk=-qj$. Since for $p=0$ the matrix $A=\mathbf{I}$, for us it will be clear that $\mathcal{M}_{0}\bullet f=f$.
@@ 1.3. Regular Blaschke products
Let us remark that $|\gamma|=1$. Using this trick and then proceeding as in the $\kappa=2$ case, one gets that the equation $B(q)=q_{0}$ is equivalent to $P_{\kappa}(q)=0$, where $P_{\kappa}$ is a slice regular polynomial of degree $\kappa$, hence the set of its roots has total multiplicity $\kappa=n+2m$, where $m$ is the number of spherical zeros and $n$ the number of non-spherical zeros (counted with their respective multiplicity).
@@ 2. The hyperbolic difference quotient and the three-points Schwarz–Pick Lemma
At a first look, since the term $(q-p)^{-\ast}$ is well defined only for $q\in\mathbb{B}$ not in the sphere $\mathbb{S}_{p}=\{q:q^{2}-q\cdot 2\operatorname{Re}p+|p|^{2}=0\}$, the domain of $\mathcal{R}_{f,p}$ is only $\mathbb{B}\setminus\mathbb{S}_{p}$. But the singularity at $\mathbb{S}_{p}$ is removable: indeed, by applying [16, Prop. 3.18] to ff(p)f-f(p), there exists a unique slice regular function $g:\mathbb{B}\longrightarrow\mathbb{B}$ for which $f(q)=f(p)+(q-p)\ast g(q)$, and it is clear that $g$ coincides with $\mathcal{R}_{f,p}$ in $\mathbb{H}\setminus\mathbb{S}_{p}$. The value in $p$ of this function coincides with the regular derivative $\partial_{C}f(p)$, whereas its value in $\overline{p}$ coincides by definition with the spherical derivative in $p$: $\partial_{S}f(p):=(2\operatorname{Im}p)^{-1}(f(p)-f(\overline{p}))$.  [^3] When $p$ is real, we will assume that the spherical derivative coincides with the Cullen derivative.
@@
(7) $f^{h}(p)=f_{p}^{\star}(p)=(1-|p|^{2})\partial_{C}f(p)(1-\overline{f(p)}f(p_{0}))^{-1},\quad\text{for some }p_{0}\in\mathbb{S}_{p}$
where $p_{0}\in\mathbb{S}_{p}$ depends on $p$ and on the function $f$ itself. Note that $f_{p}^{\star}$ can also be written as
$f_{p}^{\star}=\mathcal{M}_{p}^{-\ast}\ast(\mathcal{M}_{f(p)}\bullet f)$}
@@ 2.2. Three-points quaternionic Schwarz–Pick Lemma
3. Some consequences. Estimates in the case of $f(0)=0$
@@
Since $\mathcal{M}_{0}$ is the identity function $q\mapsto q$ and $\mathcal{M}_{0}\bullet f=f$, it is clear that $f^{\star}_{0}(q_{0})=(\mathcal{M}_{0}^{-\ast}\ast(\mathcal{M}_{0}\bullet f))(q_{0})=q_{0}^{-1}f(q_{0})$. On the other side, since 0 is invariant by the rotations appearing by evaluating the $\ast$-products punctually, then
@@
Using the fact that $\mathcal{M}_{r}=M_{r}$ is a slice preserving function and that $\mathcal{M}_{f(r)}\bullet f$ coincides with $M_{f(r)}\circ f$ on the real interval (1,1)(-1,1), one gets
@@
and the factor $(1-q\overline{p})^{-1}(1-p\overline{q})$ has clearly modulus one. ∎
@@
By the triangle inequality and (12), where we replace $f$ with $f_{r}^{\ast}$ and the two generic points $p$ and $q$ appearing there with 0 and $s$, it is
@@
3.1. Distortion estimates in terms of the hyperbolic derivative
Almost like the complex case [7], one can use the three-point Schwarz–Pick Lemma to get bounds for the values of the hyperbolic derivative of a regular self-map of the unit ball fixing the origin. As the hyperbolic derivative is related to the Cullen derivative, this estimates can be seen as distortion estimates, giving control to how the map can deform lengths and areas along slices going towards the boundary. First, we need to show how Euclidean and pseudo-hyperbolic balls are related. Indeed, they are the same but with different centers and radii: